# thanks: This paper has been submitted in conjunction with its companion Algeri et al. [1]

### v1

@@ V Summary and Discussion
Finally, we note that inconsistencies in the statistical formalism can be found in the literature concerning the statistical analysis of the SGWB. In particular, when the spherical harmonic coefficients of the SGWB sky map are estimated via maximum likelihood [cf. 23] the user implicitly assumes such quantities – as well as the corresponding sky maps– are deterministic. Therefore, they should not be simultaneously treated as zero-mean complex Gaussian random variables. A statistical analysis that enables the estimation and testing of angular power spectral models while also allowing the spherical harmonic coefficients to be random can be established. It requires, however, a revision of the modeling framework at a fundamental level. While this is beyond the scope of this manuscript, it is the subject of ongoing work.
@@ Acknowledgements.
SA and XZ are grateful for the financial support provided by the Office of the Vice President for Research & Innovation at the Uni- versity of Minnesota. XZ was partially supported by the University of Minnesota Data Science Initiative with funding made available by the MnDrive initiative. SA, GJ, and VM were partially supported by NSF grant DMS-2152746. The work of EF and VM was in part supported by NSF grants PHY-2110238 and PHY-1806630. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by NSF Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
@@ Appendix A The generalized $\chi^{2}$ distribution of the angular power spectrum
Define the index set =$\mathcal{I}_{\ell}=\{(\ell,m)\in\mathbb{N}^{+}\times\mathbb{Z};-\ell\leq m\leq\ell\}$ and define
$\mathcal{I}=\bigcup_{\ell=1}^{\ell_{max}}\mathcal{I}_{\ell}.$
@@
$\{\hat{a}_{\ell m}\}_{\mathcal{I}}\sim\mathcal{CN}(\{{a}_{\ell m}\}_{\mathcal{I}},\bm{\Gamma}^{-1}),$
@@
$\exp\{-(\hat{a}_{\ell^{\prime}m^{\prime}}-{a}_{\ell^{\prime}m^{\prime}})^{*}\Gamma_{\ell^{\prime}m^{\prime},\ell m}(\hat{a}_{\ell m}-{a}_{\ell^{\prime}m^{\prime}})\}$
$=$ $\exp\{-\tfrac{1}{2}(\bm{X}-\bm{\mu})^{T}\bm{\Omega}^{-1}(\bm{X}-\bm{\mu})\},$
@@
$\bm{X}=\begin{pmatrix}\operatorname{Re}(\{\hat{a}_{\ell m}\}_{\mathcal{I}})\\ \operatorname{Im}(\{\hat{a}_{\ell m}\}_{\mathcal{I}})\\ \end{pmatrix},\quad\bm{\mu}=\begin{pmatrix}\operatorname{Re}(\{{a}_{\ell m}\}_{\mathcal{I}})\\ \operatorname{Im}(\{{a}_{\ell m}\}_{\mathcal{I}})\\ \end{pmatrix}$
@@
$\bm{\Omega}^{-1}={2}\begin{pmatrix}\operatorname{Re}(\bm{\Gamma})&-\operatorname{Im}(\bm{\Gamma})\\ \operatorname{Im}(\bm{\Gamma})&\operatorname{Re}(\bm{\Gamma})\end{pmatrix},$
@@
$\bm{\Omega}=\frac{1}{2}\begin{pmatrix}\operatorname{Re}(\bm{\Gamma}^{-1})&-\operatorname{Im}(\bm{\Gamma}^{-1})\\ \operatorname{Im}(\bm{\Gamma}^{-1})&\operatorname{Re}(\bm{\Gamma}^{-1})\end{pmatrix}.$
@@
$\bm{X}_{\ell}=\begin{pmatrix}\operatorname{Re}(\{\hat{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \operatorname{Im}(\{\hat{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \end{pmatrix}\sim\mathcal{N}\left(\bm{\mu}_{\ell},\bm{\Omega}_{\ell}\right),$
@@
$\bm{\mu}_{\ell}=\begin{pmatrix}\operatorname{Re}(\{{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \operatorname{Im}(\{{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \end{pmatrix},$
@@
$\bm{\Omega}_{\ell}=\frac{1}{2}\begin{pmatrix}\operatorname{Re}(\bm{\Gamma}^{-1})_{\ell,\ell}&-\operatorname{Im}(\bm{\Gamma}^{-1})_{\ell,\ell}\\ \operatorname{Im}(\bm{\Gamma}^{-1})_{\ell,\ell}&\operatorname{Re}(\bm{\Gamma}^{-1})_{\ell,\ell}\end{pmatrix}.$
@@
$\bm{\Omega}_{\ell}=\bm{C}_{\ell}\bm{\Lambda}_{\ell}\bm{C}_{\ell}^{T}=\sum_{i=1}^{r_{\ell}}\lambda_{\ell,i}\bm{C}_{\ell,i}\bm{C}_{\ell,i}^{T},$
@@
$\begin{split}\bm{U}_{\ell}&=\bm{\Lambda}_{\ell}^{-1/2}\bm{C}_{\ell}^{T}(\bm{X}_{\ell}-\bm{\mu}_{\ell})=\begin{pmatrix}\bm{U}_{\ell,1}\\ \bm{U}_{\ell,2}\\ \vdots\\ \bm{U}_{\ell,r}\end{pmatrix},\end{split}$
@@
$\begin{split}\bm{\mu}_{\ell}^{\prime}&=\bm{\Lambda}_{\ell}^{-1/2}\bm{C}_{\ell}^{T}\bm{\mu}_{\ell}=\begin{pmatrix}\bm{\mu}^{\prime}_{\ell,1}\\ \bm{\mu}^{\prime}_{\ell,2}\\ \vdots\\ \bm{\mu}^{\prime}_{\ell,r}\end{pmatrix},\end{split}$
@@
$\bm{U}_{\ell}\sim\mathcal{N}(\bm{0},\bm{I}_{4\ell+2})\text{ and }\bm{X}_{\ell}=\bm{C}_{\ell}\bm{\Lambda}_{\ell}^{1/2}(\bm{U}_{\ell}+\bm{\mu}^{\prime}_{\ell}).$
@@
$\begin{split}\|\bm{X}_{\ell}\|^{2}=&\|\bm{C}_{\ell}\bm{\Lambda}_{\ell}^{1/2}(\bm{U}_{\ell}+\bm{\mu}^{\prime}_{\ell})\|^{2}\\ =&\|\bm{\Lambda}_{\ell}^{1/2}(\bm{U}_{\ell}+\bm{\mu}^{\prime}_{\ell})\|^{2}=\sum_{i=1}^{r_{\ell}}\lambda_{\ell,i}\|\bm{U}_{\ell,i}+\bm{\mu}_{\ell,i}^{\prime}\|^{2},\end{split}$
@@
$\theta_{\ell,i}=\|\bm{\mu}^{\prime}_{\ell,i}\|^{2}=\lambda_{\ell,i}^{-1}\bm{\mu}_{\ell}^{T}\bm{C}_{\ell,i}\bm{C}_{\ell,i}^{T}\bm{\mu}_{\ell}$
@@ Appendix B The covariance of the angular power spectrum
$\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}{\mu}_{U}\\ {\mu}_{V}\end{pmatrix},\begin{pmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{12}&\sigma_{22}\end{pmatrix}\right),$
@@
$\begin{split}M_{U,V}(a,b)&=\langle{e^{aU+bV}\rangle}\\ &=\exp\left(\mu_{U}a+\mu_{V}b+\frac{1}{2}a^{2}\sigma_{11}+ab\sigma_{12}+\frac{1}{2}b^{2}\sigma_{22}\right)\end{split}$
@@
$\mathbf{Z}=[Z_{1},Z_{2},\cdots,Z_{n}]^{T}\sim\mathcal{CN}(\bm{\mu},\bm{\Xi}).$
The real and imaginary parts of $\mathbf{Z}$ satisfy [Cf. 56, for details]
@@
$\begin{split}\bm{\Xi}^{\prime}&=\begin{pmatrix}\Xi_{11}&\Xi_{12}\\ \Xi_{21}&\Xi_{22}\end{pmatrix},\end{split}$
@@
$\quad\operatorname{Re}(\bm{\Xi}^{\prime})=\begin{pmatrix}\Xi_{11}&\operatorname{Re}(\Xi_{12})\\ \operatorname{Re}(\Xi_{21})&\Xi_{22}\end{pmatrix},$
@@
$\operatorname{Im}(\bm{\Xi}^{\prime})=\begin{pmatrix}0&-\operatorname{Im}(\Xi_{12})\\ \operatorname{Im}(\Xi_{21})&0\end{pmatrix}.$
@@
$\begin{split}{\rm Var}(X_{1}^{2})=&\left\langle X_{1}^{4}\right\rangle-\left\langle X_{1}^{2}\right\rangle^{2}\\ =&\frac{3}{4}\Xi_{11}^{2}+3\mu_{X_{1}}^{2}\Xi_{11}+\mu_{X_{1}}^{4}-\left(\frac{1}{2}\Xi_{11}+\mu_{X_{1}}^{2}\right)^{2}\\ =&\frac{1}{2}\Xi_{11}^{2}+2\mu^{2}_{X_{1}}\Xi_{11}\end{split}$
@@
$\{\hat{a}_{{f_{b}},s,\ell m}\}_{\mathcal{I}}\sim\mathcal{CN}(\{{a}_{{f_{b}},\ell m}\}_{\mathcal{I}},\bm{\Gamma}_{{f_{b}},s}^{-1}),$
where the index set =$\mathcal{I}=\{(\ell,m)\in\mathbb{Z}^{2};1\leq\ell\leq\ell_{max},-\ell\leq m\leq\ell\}$, and $\bm{\Gamma}_{{f_{b}},s}^{-1}=\{(\Gamma_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}\}_{(\ell,m),(\ell^{\prime},m^{\prime})\in\mathcal{I}}$, applying Lemma 2 yields
${\rm Cov}(|\hat{a}_{{f_{b}},s,\ell m}|^{2},|\hat{a}_{{f_{b}},s,\ell^{\prime}m^{\prime}}|^{2})$
@@
$(\Sigma_{{f_{b}},s})_{\ell,\ell^{\prime}}={\rm Cov}(\hat{A}_{{f_{b}},s,\ell},\hat{A}_{{f_{b}},s,\ell^{\prime}}).$
@@
$\begin{split}\sum_{m,m^{\prime}}\frac{|({\Gamma}_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}|^{2}+2\,\operatorname{Re}\left[\hat{a}^{*}_{{f_{b}},\ell m}(\Gamma_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}}\right]}{(1+2\ell)(1+2\ell^{\prime})}.\end{split}$
@@
$X_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}X.$
Let $\|\cdot\|_{op}$ denote the operator norm, which is equal to the largest singular value of the matrix.
@@
    $\max_{{f_{b}}}\max_{1\leq s\leq S}\bigl{\|}\hat{\Sigma}_{{f_{b}},s}^{-1}-\Sigma_{{f_{b}},s}^{-1}\bigr{\|}_{op}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}0.$
@@
$\{\hat{a}_{{f_{b}},\ell m}\}_{\mathcal{I}}\;\sim\;\mathcal{CN}\bigl{(}\{a_{{f_{b}},\ell m}\}_{\mathcal{I}},\,\Gamma_{{f_{b}}}^{-1}\bigr{)},$
@@
$\hat{a}_{{f_{b}},\ell m}\;\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\;a_{{f_{b}},\ell m}\quad\text{as }S\to\infty.$
@@
$\begin{split}g_{\ell,\ell^{\prime}}(\{\hat{a}_{f,\ell m}\}_{\mathcal{I}},\bm{\Omega})=\sum_{m,m^{\prime}}\frac{|{\Omega}_{\ell m,\ell^{\prime}m^{\prime}}|^{2}+2\operatorname{Re}[\hat{a}_{f,\ell m}^{*}{\Omega}_{\ell m,\ell^{\prime}m^{\prime}}\hat{a}_{f,\ell^{\prime}m^{\prime}}]}{(1+2\ell)(1+2\ell^{\prime})}.\end{split}$
By the continuous mapping theorem [Cf. 40, Ch. 2], we obtain that as $\hat{a}_{{f_{b}},\ell m}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}{a}_{{f_{b}},\ell m}$
$g_{\ell,\ell^{\prime}}(\{\hat{a}_{{f_{b}},\ell m}\}_{\mathcal{I}},\Gamma^{-1}_{{f_{b}},s})\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}g_{\ell,\ell^{\prime}}(\{{a}_{{f_{b}},\ell m}\}_{\mathcal{I}},\Gamma^{-1}_{{f_{b}},s}).$
@@
$\inf_{s\geq 1,{f_{b}}}\lambda_{min}(\Sigma_{{f_{b}},s})\geq\beta,\text{ and }\inf_{s\geq 1,{f_{b}}}\lambda_{min}(\hat{\Sigma}_{{f_{b}},s})\geq\beta,$
@@
$\hat{\Sigma}^{-1}_{{f_{b}},s}-{\Sigma}^{-1}_{{f_{b}},s}=-\hat{\Sigma}^{-1}_{{f_{b}},s}(\hat{\Sigma}_{{f_{b}},s}-{\Sigma}_{{f_{b}},s}){\Sigma}_{{f_{b}},s}^{-1},$
@@
$\begin{split}\|\hat{\Sigma}^{-1}_{{f_{b}},s}-{\Sigma}^{-1}_{{f_{b}},s}\|_{op}&\leq\|\hat{\Sigma}_{{f_{b}},s}^{-1}\|_{op}\|\hat{\Sigma}_{{f_{b}},s}-{\Sigma}_{{f_{b}},s}\|_{op}\|{\Sigma}_{{f_{b}},s}^{-1}\|_{op}\\ &\leq\frac{1}{\beta^{2}}\|\hat{\Sigma}_{{f_{b}},s}-{\Sigma}_{{f_{b}},s}\|_{op}.\end{split}$

### v3

@@ V Summary and Discussion
In an astrophysical SGWB comprised of CBC events, the temporal discreteness of the events gives rise to a shot noise which dominates the background signal [13, 57]. In the application of our statistical framework to O3 data, we do not account for this noise, but alternative estimation methods for the angular power spectrum have been developed to obtain estimates that are unbiased by temporal shot noise [57, 58]. These methods involve subdividing the data set into separate segments as we have done in this study. In our framework, we perform this subdivision to obtain multiple sets of estimators to increase our sample size, but the aforementioned unbiased methods require this subdivision to obtain only a single set of estimators. It may therefore be difficult to account for the temporal shot noise while also satisfying our statistical framework’s sample size requirements, and future work should be done to find a balance between these two factors.
Finally, we note that inconsistencies in the statistical formalism can be found in the literature concerning the statistical analysis of the SGWB. In particular, when the spherical harmonic coefficients of the SGWB sky map are estimated via maximum likelihood [cf. 25] the user implicitly assumes such quantities – as well as the corresponding sky maps– are deterministic. Therefore, they should not be simultaneously treated as zero-mean complex Gaussian random variables. A statistical analysis that enables the estimation and testing of angular power spectral models while also allowing the spherical harmonic coefficients to be random can be established. It requires, however, a revision of the modeling framework at a fundamental level. While this is beyond the scope of this manuscript, it is the subject of ongoing work.
@@ Acknowledgements.
SA and XZ are grateful for the financial support provided by the Office of the Vice President for Research & Innovation at the University of Minnesota. XZ was partially supported by the University of Minnesota Data Science Initiative with funding made available by the MnDrive initiative. SA, GJ, and VM were partially supported by NSF grant DMS-2152746. The work of EF and VM was in part supported by NSF grants PHY-2110238 and PHY-1806630. HZ was supported by the IRSA Faragher Distinguished Postdoctoral Fellowship. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by NSF Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.
@@ Appendix A The generalized $\chi^{2}$ distribution of the angular power spectrum
Define the index set $\mathcal{I}_{\ell}=\{(\ell,m)\in\mathbb{N}^{+}\times\mathbb{Z};-\ell\leq m\leq\ell\}$ and define
$\mathcal{I}=\bigcup_{\ell=1}^{\ell_{\rm max}}\mathcal{I}_{\ell}.$ (75)
@@
$\{\hat{a}_{\ell m}\}_{\mathcal{I}}\sim\mathcal{CN}(\{{a}_{\ell m}\}_{\mathcal{I}},\bm{\Gamma}^{-1}),$ (76)
@@
$\exp\{-(\hat{a}_{\ell^{\prime}m^{\prime}}-{a}_{\ell^{\prime}m^{\prime}})^{*}\Gamma_{\ell^{\prime}m^{\prime},\ell m}(\hat{a}_{\ell m}-{a}_{\ell^{\prime}m^{\prime}})\}$ (78)
$=$ $\exp\{-\tfrac{1}{2}(\bm{X}-\bm{\mu})^{T}\bm{\Omega}^{-1}(\bm{X}-\bm{\mu})\},$ (79)
@@
$\bm{X}=\begin{pmatrix}\operatorname{Re}(\{\hat{a}_{\ell m}\}_{\mathcal{I}})\\ \operatorname{Im}(\{\hat{a}_{\ell m}\}_{\mathcal{I}})\\ \end{pmatrix},\quad\bm{\mu}=\begin{pmatrix}\operatorname{Re}(\{{a}_{\ell m}\}_{\mathcal{I}})\\ \operatorname{Im}(\{{a}_{\ell m}\}_{\mathcal{I}})\\ \end{pmatrix}$ (80)
@@
$\bm{\Omega}^{-1}={2}\begin{pmatrix}\operatorname{Re}(\bm{\Gamma})&-\operatorname{Im}(\bm{\Gamma})\\ \operatorname{Im}(\bm{\Gamma})&\operatorname{Re}(\bm{\Gamma})\end{pmatrix}$, (81)
@@
$\bm{\Omega}=\frac{1}{2}\begin{pmatrix}\operatorname{Re}(\bm{\Gamma}^{-1})&-\operatorname{Im}(\bm{\Gamma}^{-1})\\ \operatorname{Im}(\bm{\Gamma}^{-1})&\operatorname{Re}(\bm{\Gamma}^{-1})\end{pmatrix}$. (82)
@@
$\bm{X}_{\ell}=\begin{pmatrix}\operatorname{Re}(\{\hat{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \operatorname{Im}(\{\hat{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \end{pmatrix}\sim\mathcal{N}\left(\bm{\mu}_{\ell},\bm{\Omega}_{\ell}\right),$ (83)
@@
$\bm{\mu}_{\ell}=\begin{pmatrix}\operatorname{Re}(\{{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \operatorname{Im}(\{{a}_{\ell m}\}_{\mathcal{I}_{\ell}})\\ \end{pmatrix}$, (84)
@@
$\bm{\Omega}_{\ell}=\frac{1}{2}\begin{pmatrix}\operatorname{Re}(\bm{\Gamma}^{-1})_{\ell,\ell}&-\operatorname{Im}(\bm{\Gamma}^{-1})_{\ell,\ell}\\ \operatorname{Im}(\bm{\Gamma}^{-1})_{\ell,\ell}&\operatorname{Re}(\bm{\Gamma}^{-1})_{\ell,\ell}\end{pmatrix}$. (85)
@@
$\bm{\Omega}_{\ell}=\bm{C}_{\ell}\bm{\Lambda}_{\ell}\bm{C}_{\ell}^{T}=\sum_{i=1}^{r_{\ell}}\lambda_{\ell,i}\bm{C}_{\ell,i}\bm{C}_{\ell,i}^{T},$ (87)
@@
[…]
@@
[…]
@@
$\bm{U}_{\ell}\sim\mathcal{N}(\bm{0},\bm{I}_{4\ell+2})\text{ and }\bm{X}_{\ell}=\bm{C}_{\ell}\bm{\Lambda}_{\ell}^{1/2}(\bm{U}_{\ell}+\bm{\mu}^{\prime}_{\ell}).$ (90)
@@
$\begin{split}\|\bm{X}_{\ell}\|^{2}=&\|\bm{C}_{\ell}\bm{\Lambda}_{\ell}^{1/2}(\bm{U}_{\ell}+\bm{\mu}^{\prime}_{\ell})\|^{2}\\ =&\|\bm{\Lambda}_{\ell}^{1/2}(\bm{U}_{\ell}+\bm{\mu}^{\prime}_{\ell})\|^{2}=\sum_{i=1}^{r_{\ell}}\lambda_{\ell,i}\|\bm{U}_{\ell,i}+\bm{\mu}_{\ell,i}^{\prime}\|^{2},\end{split}$ (91)
@@
$\theta_{\ell,i}=\|\bm{\mu}^{\prime}_{\ell,i}\|^{2}=\lambda_{\ell,i}^{-1}\bm{\mu}_{\ell}^{T}\bm{C}_{\ell,i}\bm{C}_{\ell,i}^{T}\bm{\mu}_{\ell}$ (92)
@@ Appendix B The covariance of the angular power spectrum
$\begin{pmatrix}U\\ V\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}{\mu}_{U}\\ {\mu}_{V}\end{pmatrix},\begin{pmatrix}\sigma_{11}&\sigma_{12}\\ \sigma_{12}&\sigma_{22}\end{pmatrix}\right)$, (95)
@@
$\begin{split}M_{U,V}(a,b)&=\langle{e^{aU+bV}\rangle}\\ &=\exp\left(\mu_{U}a+\mu_{V}b+\frac{1}{2}a^{2}\sigma_{11}+ab\sigma_{12}+\frac{1}{2}b^{2}\sigma_{22}\right)\end{split}$ (98)
@@
$\mathbf{Z}=[Z_{1},Z_{2},\cdots,Z_{n}]^{T}\sim\mathcal{CN}(\bm{\mu},\bm{\Xi}).$ (100)
The real and imaginary parts of $\mathbf{Z}$ satisfy [Cf. 65, for details]
@@
[…]
@@
$\quad\operatorname{Re}(\bm{\Xi}^{\prime})=\begin{pmatrix}\Xi_{11}&\operatorname{Re}(\Xi_{12})\\ \operatorname{Re}(\Xi_{21})&\Xi_{22}\end{pmatrix}$, (105)
@@
$\operatorname{Im}(\bm{\Xi}^{\prime})=\begin{pmatrix}0&-\operatorname{Im}(\Xi_{12})\\ \operatorname{Im}(\Xi_{21})&0\end{pmatrix}$. (106)
@@
$\begin{split}{\rm Var}(X_{1}^{2})=&\left\langle X_{1}^{4}\right\rangle-\left\langle X_{1}^{2}\right\rangle^{2}\\ =&\frac{3}{4}\Xi_{11}^{2}+3\mu_{X_{1}}^{2}\Xi_{11}+\mu_{X_{1}}^{4}-\left(\frac{1}{2}\Xi_{11}+\mu_{X_{1}}^{2}\right)^{2}\\ =&\frac{1}{2}\Xi_{11}^{2}+2\mu^{2}_{X_{1}}\Xi_{11}\end{split}$ (112)
@@
$\{\hat{a}_{{f_{b}},s,\ell m}\}_{\mathcal{I}}\sim\mathcal{CN}(\{{a}_{{f_{b}},\ell m}\}_{\mathcal{I}},\bm{\Gamma}_{{f_{b}},s}^{-1}),$ (115)
where the index set $\mathcal{I}=\{(\ell,m)\in\mathbb{Z}^{2};1\leq\ell\leq\ell_{\rm max},-\ell\leq m\leq\ell\}$, and $\bm{\Gamma}_{{f_{b}},s}^{-1}=\{(\Gamma_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}\}_{(\ell,m),(\ell^{\prime},m^{\prime})\in\mathcal{I}}$, applying Lemma 2 yields
${\rm Cov}(|\hat{a}_{{f_{b}},s,\ell m}|^{2},|\hat{a}_{{f_{b}},s,\ell^{\prime}m^{\prime}}|^{2})$ (116)
@@
$(\Sigma_{{f_{b}},s})_{\ell,\ell^{\prime}}={\rm Cov}(\hat{A}_{{f_{b}},s,\ell},\hat{A}_{{f_{b}},s,\ell^{\prime}}).$ (118)
@@
$\begin{split}\sum_{m,m^{\prime}}\frac{|({\Gamma}_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}|^{2}+2\,\operatorname{Re}\left[\hat{a}^{*}_{{f_{b}},\ell m}(\Gamma_{{f_{b}},s}^{-1})_{\ell m,\ell^{\prime}m^{\prime}}\hat{a}_{{f_{b}},\ell^{\prime}m^{\prime}}\right]}{(1+2\ell)(1+2\ell^{\prime})}.\end{split}$ (119)
@@
$X_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}X.$ (120)
Let $\|\cdot\|_{\rm op}$ denote the operator norm, which is equal to the largest singular value of the matrix.
@@
    $\max_{{f_{b}}}\max_{1\leq s\leq S}\bigl\|\hat{\Sigma}_{{f_{b}},s}^{-1}-\Sigma_{{f_{b}},s}^{-1}\bigr\|_{\rm op}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}0.$ (122)
@@
$\{\hat{a}_{{f_{b}},\ell m}\}_{\mathcal{I}}\;\sim\;\mathcal{CN}\bigl(\{a_{{f_{b}},\ell m}\}_{\mathcal{I}},\,\Gamma_{{f_{b}}}^{-1}\bigr),$ (125)
@@
$\hat{a}_{{f_{b}},\ell m}\;\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\;a_{{f_{b}},\ell m}\quad\text{as }S\to\infty.$ (126)
@@
$\begin{split}g_{\ell,\ell^{\prime}}(\{\hat{a}_{f,\ell m}\}_{\mathcal{I}},\bm{\Omega})=\sum_{m,m^{\prime}}\frac{|{\Omega}_{\ell m,\ell^{\prime}m^{\prime}}|^{2}+2\operatorname{Re}[\hat{a}_{f,\ell m}^{*}{\Omega}_{\ell m,\ell^{\prime}m^{\prime}}\hat{a}_{f,\ell^{\prime}m^{\prime}}]}{(1+2\ell)(1+2\ell^{\prime})}.\end{split}$ (127)
By the continuous mapping theorem [Cf. 46, Ch. 2], we obtain that as $\hat{a}_{{f_{b}},\ell m}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}{a}_{{f_{b}},\ell m}$
$g_{\ell,\ell^{\prime}}(\{\hat{a}_{{f_{b}},\ell m}\}_{\mathcal{I}},\Gamma^{-1}_{{f_{b}},s})\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}g_{\ell,\ell^{\prime}}(\{{a}_{{f_{b}},\ell m}\}_{\mathcal{I}},\Gamma^{-1}_{{f_{b}},s}).$ (128)
@@
[…]
@@
$\hat{\Sigma}^{-1}_{{f_{b}},s}-{\Sigma}^{-1}_{{f_{b}},s}=-\hat{\Sigma}^{-1}_{{f_{b}},s}(\hat{\Sigma}_{{f_{b}},s}-{\Sigma}_{{f_{b}},s}){\Sigma}_{{f_{b}},s}^{-1},$ (130)
@@
$\begin{split}\|\hat{\Sigma}^{-1}_{{f_{b}},s}-{\Sigma}^{-1}_{{f_{b}},s}\|_{\rm op}&\leq\|\hat{\Sigma}_{{f_{b}},s}^{-1}\|_{\rm op}\|\hat{\Sigma}_{{f_{b}},s}-{\Sigma}_{{f_{b}},s}\|_{\rm op}\|{\Sigma}_{{f_{b}},s}^{-1}\|_{\rm op}\\ &\leq\frac{1}{\beta^{2}}\|\hat{\Sigma}_{{f_{b}},s}-{\Sigma}_{{f_{b}},s}\|_{\rm op}.\end{split}$ (131)